OpenStudy (xmoses1):

If the point (-1/2, y) lies on the line whose equation is 2x - 3y = 6, what is the value of y? -7/3 -5/3 7/3

OpenStudy (xmoses1):

@jim_thompson5910 could you help me out?

jimthompson5910 (jim_thompson5910):

the point (-1/2, y) means that x = -1/2 and y is currently unknown

jimthompson5910 (jim_thompson5910):

2x - 3y = 6 2(-1/2) - 3y = 6 ... replace x with -1/2 (since x = -1/2) -2/2 - 3y = 6 -1 - 3y = 6 I'll let you finish up

OpenStudy (xmoses1):

Ok so -3y=7?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (xmoses1):

so -3/7

OpenStudy (xmoses1):

oops i mean -7/3 lol

jimthompson5910 (jim_thompson5910):

yep, -7/3 is the answer

OpenStudy (xmoses1):

Ok so all i had to do was insert the number that was given to me

OpenStudy (xmoses1):

Thanks again :)

jimthompson5910 (jim_thompson5910):

you're welcome